The product of two numbers is 90.
One number is 3 more than twice the other number.
Model these numbers with a quadratic equation.
Solve the equation by factoring and using the Zero-Product Property.

Answer by ewatrrr(24785)

(2n+3)n = 90
2n^2 + 3n - 90 = 0
(2n+15)(n-6) = 0
n = -15/2, 6